# Galois theory, automorphism groups of prime models, and the Picard-Vessiot closure

### v1

@@ Abstract
The main result of the paper extends the well-known Galois correspondence between closed subgroups of $Aut(acl^{eq}(A)/A)$ and dcldcl-closed intermediate sets $A\subseteq C\subseteq acl^{eq}(A)$ to the setting of the minimal closure inside of a prime model of a totally transcendental theory $T=T^{eq}$, $A\subseteq mcl^{eq}(A)\subseteq M_{A}$ (namely, $mcl^{eq}(A)$ is the intersection of all self elementary embeddings of the prime model $M_{A}$). Precisely, for $B$ a normal intermediate extension $A\subseteq B\subseteq mcl^{eq}(A)$, definable subgroups of $Aut(B/A)$ are in Galois correspondence with dcldcl-closed intermediate sets $A\subseteq C\subseteq B$.
This applies in particular to the “Picard-Vessiot closure” $K^{PV_{\infty}}$ (or $K_{\infty}$) of a differential field $K$ of characteristic 0 with algebraically closed field $C_{K}$ of constants. We also show that normal differential subfields of $K^{PV_{\infty}}$ containing $K$ are “iterated $PV$-extensions” of $K$, and the Galois correspondence result above holds for these extensions. This fills in some missing parts of Magid’s paper [5].
@@ 1 Introduction
In this paper we work in the context of a totally transcendental theory $T$ eliminating imaginaries ($T=T^{eq}$). $\bar{M}$ will denote a very saturated model, which exists as $T$ is totally transcendental ($t.t.$). Typically, $A$ and $B$ will denote small definably closed subsets of $\bar{M}$. For $A\subseteq B$, $Aut(B/A)$ denotes the group of $A$-elementary permutations of $B$. In the context of a $t.t.$ theory, one has, over arbitrary parameter sets $A$, prime models $M_{A}$ which are unique up to $A$-isomorphism and which are atomic and strongly $\omega$-homogeneous over $A$ (whenever $\bar{a}$, and $\bar{b}$ are finite tuples in $M_{A}$ with the same type over $A$, there is an automorphism $\sigma\in Aut(M_{A}/A)$ taking $\bar{a}$ to $\bar{b}$).
@@
There are two main notions used in this paper to set up the Galois correspondence. Firstly, is that of a (relatively) definable subset of an automorphism group of a prime model. These were defined in [9] and used to show existence and uniqueness of invariant measures on automorphism groups under certain assumptions. In [9], this notion was used to give a characterization of definable cocycles (first defined in [8]). This was extended in [7]. Also in [9] this notion is related to differential Galois theory.
The second notion used to set up the correspondence is the minimal closure of a set of parameters $A$ contained in the prime model $M_{A}$, which we denote $mcl(A)$. The minimal closure $mcl(A)$ is the intersection of the images all elementary embeddings of the prime model $M_{A}$ into itself.
In [13], $mcl(A)$ is defined and strong homogeneity and closure properties of $mcl(A)$ are shown in anticipation of [12] where they were used to give the first model-theoretic description of the Galois correspondence for the Picard-Vessiot, and strongly normal differential Galois theories. The other major result of [12] is the well-known “infinite” Galois correspondence for aclacl, developed using the profinite topology on the automorphism group. However, in [12], working in $DCF_{0}$, the theory of differential fields of characteristic 0, (essentially the context of a prime model of a $t.t.$ theory) only the “finite type” Picard-Vessiot extensions are studied. No topology or definable structure is given to the automorphism groups of infinitely dcldcl-generated extensions. The structure on the automorphism groups associated to finitely dcldcl-generated extensions is given by the binding group construction.
In this paper, we replace the role of the topology of pointwise convergence (or the profinite topology) on $Aut(acl^{eq}(A)/A)$ with the notion from [9] of a definable group of automorphisms to give a satisfying “infinite” Galois theory for $Aut(mcl^{eq}(A)/A)$.
Before stating this main theorem we give a few more definitions and say something about our approach. Let $A$ be a parameter set and $M_{A}$ be a copy of the prime model over $A$. Then an intermediate set $A\subseteq B\subseteq M_{A}$ is said to be normal (in $M_{A}$ over $A$) if $B$ is fixed setwise by $Aut(M_{A}/A)$ in which case $Aut(M_{A}/A)$ has $Aut(M_{A}/B)$ as a normal subgroup with a surjective quotient map to $Aut(B/A)$ given by restriction. Next, we call $B$ minimal (in $M_{A}$ over $A$) if there is no $C$ with $A\subseteq C\subseteq B$ such that $C$ is Tarski-Vaught in $B$. We show that, $mcl(A)$, is the largest (under inclusion) normal minimal subset of $M_{A}$ over $A$, i.e. any normal minimal subset of $M_{A}$ over $A$ is contained in the minimal closure and of course any normal subset $B$ of $mcl(A)$ containing $A$ must be minimal.
The main result is:
@@
For $B$, a normal and minimal set in $M_{A}$ over $A$, equivalently an intermediate extension $A\subseteq B\subseteq mcl^{eq}(A)$, definable subgroups of $Aut(B/A)$ correspond to dcldcl-closed intermediate sets $A\subseteq C\subseteq B$.
In fact, our statement in Section 2 (Theorem 2.13) will state this correspondence for normal minimal sets without explicit reference to the minimal closure; the reason being that we find it more comfortable to work with the notion of Tarski-Vaught minimality than with the minimal closure, although we end section 2 with the necessary comparison of notions.
@@
Already in [12] and [13] where the minimal closure is introduced, it is stated that $K^{PV}$ is normal and minimal being contained in the minimal closure $K\subseteq K^{PV}\subseteq mcl(K)\subseteq K^{diff}$. But in fact, more is true:
@@
Each of $K^{PV_{n}}$ and $K^{PV_{\infty}}$ are normal and minimal over $K$ in $K^{diff}$, i.e. $K\subseteq K^{PV}\subseteq(K^{PV})^{PV}\subseteq...\subseteq K^{PV_{\infty}}\subseteq mcl(K)\subseteq K^{diff}$.
@@ 2 The Galois theory
Finally in this section we make the connection with works of Poizat referred to above. In the context of $T$ totally transcendental, Poizat defines $A_{min}$, the minimal closure of $A$ to be the intersection of all $f(M_{A})$ as $f$ ranges over $A$-elementary embeddings of $M_{A}$ into itself. The notation $A_{min}$ comes from [11]. In [13] Poizat uses the notation $Cl(A)$ for $A_{min}$ where an equivalent definition of $A_{min}$ is given, in terms of atomic algebraicity. In this paper we will use the notation $mcl(A)$.
@@
    $mcl(A)$ is the greatest (under inclusion) $B$ with $A\subseteq B\subseteq M_{A}$ such that $B$ is normal and minimal (over $A$ in $M_{A}$).
@@
    Any subset of $mcl(A)$ is minimal.
@@
    Corollaire 7 of [13] (actually there are two Corollaire 7’s) say that $f(mcl(A))=mcl(A)$ for every $A$-automomorphism of $M_{A}$, so we have normality. Corollaire 5 of [13] says in particular that $mcl(A)$ is atomic over any subset of itself which contains $A$, so if $B\supseteq A$ is a proper subset of $mcl(A)$, we could not have that $B\leq_{TV}mcl(A)$.
@@
    As $B$ is normal, $f(B)\subseteq B$ and we clearly have $f(B)\leq_{TV}B$. By minimality of $B$, $f(B)=B$. Hence $B\subseteq f(M)$, so $B\subseteq mcl(A)$.
@@
    Let $B$ be a subset of $mcl(A)$. Suppose $B$ is not minimal. Then there is a $C\leq_{TV}B$ with $C$ properly contained in $B$. Let $b\in B\backslash C$, then as $M_{A}$ is still prime over any subset of $mcl(A)$ [11], $M_{A}$ is prime over $C$. So let $\phi(\bar{x})$ isolate $tp(b/C)$. By $C\leq_{TV}B$, there is a $c\in C$ realizing $\phi(\bar{x})$, so the formula $x=c$ is in this type. This contradicts $b\in B\backslash C$.
@@ 3 The Picard-Vessiot closure
(i) Each of $K^{PV_{n}}$ and $K^{PV_{\infty}}$ is normal and minimal over $K$ inside $K^{diff}$, i.e. $K\subseteq K^{PV}\subseteq(K^{PV})^{PV}\subseteq...\subseteq K^{PV_{\infty}}\subseteq mcl(K)\subseteq K^{diff}$.

### v2

@@ Abstract
We work in the context of a complete totally transcendental theory $T=T^{eq}$. We consider the prime model $M_{A}$ over a set $A$. For intermediate sets $B$ with $A\subseteq B\subseteq M_{A}$ which are normal ($Aut(M_{A}/A)$-invariant) and “minimal” we give a full Galois correspondence between intermediate definably closed sets $A\subseteq B\subseteq M_{A}$ and “closed” subgroups of $Aut(B/A)$ (the group of $A$-elementary permutations of $B$).
The unique greatest such minimal normal $B$ coincides with Poizat’s “minimal closure” $A_{min}$, so our paper can be see as extending the well-known Galois correspondence between closed subgroups of the profinite group $Aut(acl(A)/A)$ and intermediate definably closed sets, to the case of $A_{min}$ in place of $acl(A)$.
The main result applies in particular to the “Picard-Vessiot closure” $K^{PV_{\infty}}$ (or $K_{\infty}$) of a differential field $K$ of characteristic 0 with algebraically closed field $C_{K}$ of constants. We also show that normal differential subfields of $K^{PV_{\infty}}$ containing $K$ are “iterated $PV$-extensions” of $K$, and the Galois correspondence result above holds for these extensions. This fills in some missing parts of Magid’s paper [5].
@@ 1 Introduction
This paper extends and builds on the paper [9] which introduced (relatively) definable subsets of automorphism groups of prime models, proved the existence and uniqueness of invariant measures on these automorphism groups, and related these notions to differential Galois theory and definable Galois cohomology. In particular the notion was used to characterize definable cocycles (first defined in [8], and was extended in [7]).
The current paper is closely related to work of Poizat (see [12], [13] and Section 2, Chapter 18 of [11]) and can even be thought of as a continuation of it. Sections 3 and 4 of the current paper are also closely related to the unpublished paper [5] of Andy Magid, which in several senses we complete and extend.
We are concerned here with a full Galois correspondence. In [9] we worked mainly in the context of a stable theory $T$, and $M$ a model of $T$ which is atomic and strongly $\omega$-homogeneous over a subset $A$. (Strongly $\omega$-homogeneous over $A$ means that whenever $\bar{a}$}, $\bar{b}$} are finite tuples in $M$ with the same type over $A$ then there is an automorphism $f$ of $M$ fixing $A$ pointwise and taking $\bar{a}$} to $\bar{b}$}.) In this paper it is convenient to assume that $T$ is totally transcendental (called $\omega$-stable when $T$ is countable.) This guarantees the existence and uniqueness of prime models over arbitrary subsets which in particular have the properties above (atomicity and strong $\omega$-homogeneity) and others too. So from now on $T$ will denote a complete, totally transcendental theory.
We will be working freely in $T^{eq}$ and let $\bar{M}$} denote a very saturated model (which exists as $T$ is $t.t$). $A,B$ will typically denote small definably closed subsets of $\bar{M}$}. For $A\subseteq B$, $Aut(B/A)$ denotes the group of $A$-elementary permutations of $B$.
We let $M_{A}$ denote the prime model over $A$ (which is unique up to $A$ isomorphism). Before stating the results we give a quick account of the main notions, which will be elaborated on later. Let $A\subseteq B\subseteq M_{A}$. We call $B$ invariant or normal (over $A$ in $M$) if $B$ is fixed setwise by $Aut(M_{A}/A)$. We call $B$ minimal (over $A$ in $M$) if there is no proper subset $C$ of $B$ with $A\subseteq C$ and $C$ Tarski-Vaught in $B$. Let $\bar{b}$} be an enumeration of $B$, $p_{0}({\bar{x}})=tp({\bar{b}}/A)$ and $S_{p_{0}}(B)$ the (Stone) space of extensions of $p_{0}$ to complete types $q({\bar{x}})$ over $B$.
For $X$ a closed subset of $S_{p_{0}}(B)$ given by a set $\Sigma({\bar{x}},{\bar{b}})$ of formulas, by $X(Aut(B/A))$ we mean $\{\sigma\in Aut(B/A):\models\Sigma(\sigma({\bar{b}}),{\bar{b}})\}$, and $X(Aut({\bar{M}}/A))=\{\sigma\in Aut({\bar{M}}/A):\models\Sigma(\sigma({\bar{b}}),{\bar{b}})\}$. When $X$ is clopen, then $X$ is determined by $X(Aut(B/A))$ as pointed out in [9], but not necessarily when $X$ is arbitrary closed. On the other hand $X$ is determined by $X(Aut({\bar{M}}/A))$. By a closed subgroup of $Aut(B/A)$ we mean a subgroup of the form $X(Aut(B/A))$ where $X$ is a closed subset of $S_{p_{0}}(B)$ and $X(Aut({\bar{M}}/A))$ is a subgroup of $Aut({\bar{M}}/A)$.
The main result proved in Section 2 is:
@@
Suppose that $A\subseteq B\subseteq M_{A}$ where $B$ is normal (over $A$ in $M_{A})$ and minimal. Then there is a Galois correspondence between closed subgroups of $Aut(B/A)$ and sets $C$ with $A\subseteq C\subseteq B$.
When $B=acl(A)$ (without the need for any $t.t.$ assumption on $T$), this is Theorem 14 of [12], after noting as in [9] that closed subgroups of $Aut(acl(A)/A)$ in the sense of this paper coincide with closed subgroups of $Aut(acl(A)/A)$ as a profinite group. Also at the end of Section 2, we make the connection with Poizat’s minimal closure from [13].
@@
Already in [12] and [13] where the minimal closure is introduced, it is stated that $K^{PV}$ is normal and minimal being contained in the minimal closure $K_{min}$ of $K$ (see also Lemma 2.14 of the current paper). But more is true.
@@
Each of $K^{PV_{n}}$ and $K^{PV_{\infty}}$ are normal and minimal over $K$ in $K^{diff}$, (so by Lemma 2.14 contained in $K_{min}$).
@@ 2 The Galois theory
Finally in this section we make the connection with works of Poizat referred to above. In the context of $T$ totally transcendental, Poizat defines $A_{min}$, the minimal closure of $A$ to be the intersection of all $f(M_{A})$ as $f$ ranges over $A$-elementary embeddings of $M_{A}$ into itself. The notation $A_{min}$ comes from [11]. In [13] Poizat uses the notation $Cl(A)$ for $A_{min}$ where an equivalent definition of $A_{min}$ is given, in terms of atomic algebraicity.
@@
    $A_{min}$ is the greatest (under inclusion) $B$ with $A\subseteq B\subseteq M_{A}$ such that $B$ is normal and minimal (over $A$ in $M_{A}$).
@@
    Any subset $B$ of $A_{min}$ is minimal (namely there is no proper subset $C$ of $B$ containing $A$ with $C\leq_{TV}B$).
@@
    Corollaire 7 of [13] (actually there are two Corollaire 7’s) say that $f(A_{min})=A_{min}$ for every $A$-automomorphism of $M_{A}$, so we have normality. Corollaire 5 of [13] says in particular that $A_{min}$ is atomic over any subset of itself which contains $A$, so if $B\supseteq A$ is a proper subset of $A_{min}$, we could not have that $B\leq_{TV}A_{min}$.
@@
    As $B$ is normal, $f(B)\subseteq B$ and we clearly have $f(B)\leq_{TV}B$. By minimality of $B$, $f(B)=B$. Hence $B\subseteq f(M)$, so $B\subseteq A_{min}$.
@@
    By part 1 and Lemma 2.11 (or Corollaire 5 of [13]), $M_{A}$ is prime over any subset $C$ of itself which contains $A$. The desired result follows.
@@ 3 The Picard-Vessiot closure
(i) Each of $K^{PV_{n}}$ and $K^{PV_{\infty}}$ is normal and minimal over $K$ inside $K^{diff}$ (so by Lemma 2.14 contained in $K_{min}$).